# On linearisation and uniqueness of preduals

### v2

@@ 4. Linearisation of vector-valued functions
Hence, by the $\beta$-boundedness of $B$ there is some $C\geq 0$ such that $|e^{\prime}(A(x^{\prime}))|\leq C$ for all $x^{\prime}\in B\cap Y_{U}$ and $e^{\prime}\in G$. We deduce that the set $T(B\cap Y_{U})$ is $\sigma(E,G)$-bounded and thus bounded in $E$ because $G$ determines boundedness. We observe that $(\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime},\beta)$ is a Fréchet space by Proposition 3.5 (b) and so its linear subspace $Y_{U}$ is metrisable. We conclude that $A\colon(Y_{U},\beta_{\mid Y_{U}})\to E$ is continuous by [27, Proposition 24.10, p. 282] as the metrisable space $(Y_{U},\beta_{\mid Y_{U}})$ is bornological by [27, Proposition 24.13, p. 283]. Since $Y_{U}$ is $\beta$-dense in $\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime}$, there is a unique continuous linear extension $\widetilde{A}\colon\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime}\to E$ of $A$ by [18, 3.4.2 Theorem, p. 61–62]. Setting $F\coloneqq\chi\circ\widetilde{A}\in\mathcal{F}(\Omega,E)_{\sigma}$ by Corollary 3.6 (a) and Theorem 4.5, we observe that
@@
In particular, Theorem 4.8 is applicable to the spaces $\mathcal{F}(\Omega)=\mathcal{VC}_{P}(\Omega)$ from Example 3.7 (ii) by [25, Corollary 4.7, p. 20–21].
@@ 5. Equivalence and uniqueness of preduals
Moreover, (the restriction of) $T_{Z}(\widetilde{\delta})$ is a bijective map from the span of $\{\delta(x)\;|\;x\in\Omega\}$ to the span of $\{\widetilde{\delta}(x)\;|\;x\in\Omega\}$, which is dense in $Z$ by [25, Proposition 2.7, p. 6], because $T_{Z}(\widetilde{\delta})(\delta(x))=\widetilde{\delta}(x)$ for all $x\in\Omega$. Thus $T_{Z}(\widetilde{\delta})$ has dense range.
@@
Let $z\in Z$. Since the span of $\{\widetilde{\delta}(x)\;|\;x\in\Omega\}$ is dense in $Z$ by [25, Proposition 2.7, p. 6], there is a net $(z_{\iota})_{\iota\in I}$ converging to $z$ and where all $z_{\iota}$ can be represented as $z_{\iota}=\sum_{x\in\Omega}a_{x,\iota}\widetilde{\delta}(x)$ with finitely many non-zero $a_{x,\iota}\in\mathbb{K}$. Using that $T_{Z}(\widetilde{\delta})^{\prime\prime}\circ j_{Y}=j_{Z}\circ T_{Z}(\widetilde{\delta})$ and setting $y_{\iota}\coloneqq\sum_{x\in\Omega}a_{x,\iota}\delta(x)\in Y$ for $\iota\in I$, we get
@@
Proposition 5.1 (d) combined with Proposition 5.2 improves [1, Proposition 3.1 (c), p. 11], [16, Theorem 3.1 (Linearization Theorem), p. 128], [29, Theorem 2.1, p. 869] and [33, Theorem 3.5, p. 19] since it shows that the Banach space $Y$ of a strong linearisation $(\delta,Y,T)$ of a Banach space $\mathcal{F}(\Omega)$ such that $T$ is an isometry is uniquely determined up to an isometric isomorphism without the need of existence of isometric isomorphisms $\widetilde{T}_{E}\colon\mathcal{F}(\Omega,E)_{\sigma}\to L(Z,E)$ such that $\widetilde{T}_{E}(f)\circ\widetilde{\delta}=f$ for all $f\in\mathcal{F}(\Omega)$ for every Banach space $E$ over $\mathbb{K}$. Further, Proposition 5.1 (d) combined with Proposition 5.2 also implies the corresponding result for the (completion of the) projective tensor product of two Banach spaces given in [35, Proposition 1.5 (Uniqueness of the Tensor Product), p. 6] and [35, Theorem 2.9, p. 22].
@@
Then it holds (a)normal-$\Rightarrow$(b). If $\mathcal{F}(\Omega)$ is a BC-space such that $\delta_{x}\in\mathcal{F}(\Omega)^{\prime}$ for all $x\in\Omega$, then it holds (b)normal-$\Rightarrow$(a).
@@
In particular, $(\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime},\beta)$ is a complete DF-space (see [25, Corollary 3.23, p. 16]) and thus webbed by [18, 12.4.6 Proposition, p. 260] if $(\mathcal{F}(\Omega),\tau)$ is a Fréchet space.
@@
Using Proposition 3.5 (c) and Corollary 3.6 (a), we get the following corollary for completely normable spaces $\mathcal{F}(\Omega)$.
@@
    It holds $(Z,\widetilde{T})\sim((\mathcal{F}(\Omega),\gamma)^{\prime},\mathcal{I})$.
@@
Now, we note that $T(\widetilde{\delta})\circ\widetilde{T}(\delta)\in L(Z)$, $\widetilde{T}(\delta)\circ T(\widetilde{\delta})\in L(Y)$ and $(T(\widetilde{\delta})\circ\widetilde{T}(\delta))(\widetilde{\delta}(x))=T(\widetilde{\delta})(\delta(x))=\widetilde{\delta}(x)$ as well as $(\widetilde{T}(\delta)\circ T(\widetilde{\delta}))(\delta(x))=\widetilde{T}(\delta)(\widetilde{\delta}(x))=\delta(x)$ for all $x\in\Omega$. Since the span of $\{\widetilde{\delta}(x)\;|\;x\in\Omega\}$ is dense in $Z$ and the span of $\{\delta(x)\;|\;x\in\Omega\}$ dense in $Y$ by [25, Proposition 2.7, p. 6], we obtain that $T(\widetilde{\delta})$ and $\widetilde{T}(\delta)$ are topological isomorphisms with $T(\widetilde{\delta})=\widetilde{T}(\delta)^{-1}$.
@@
$\widetilde{T}_{\mathbb{K}}=T_{\mathbb{K}}\circ(T_{Z}(\widetilde{\delta})^{t})^{-1}=T_{\mathbb{K}}\circ(T_{Z}(\widetilde{\delta})^{-1})^{t}$
@@
(b)$\Rightarrow$(a) Let $(Z,\varphi)$ be a predual of $(\mathcal{F}(\Omega),\tau)$ such that $Z\in\mathcal{C}$. Due to [25, Proposition 3.21 (a), p. 15] $(\mathcal{F}(\Omega),\tau)$ satisfies $\operatorname{(BBC)}$ and $\operatorname{(CNC)}$ for $\sigma_{\varphi}(\mathcal{F}(\Omega),Z)$ because $Z$ is complete and barrelled. By assumption for every $x\in\Omega$ there is $z_{x}\in Z$ with $\delta_{x}=\varphi(\cdot)(z_{x})$ and thus $\delta_{x}\in\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime}$ by [25, Proposition 3.21 (b), (c), p. 15]. Hence $(\Delta,\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})$ is a strong complete barrelled linearisation of $\mathcal{F}(\Omega)$ by Corollary 3.6 (a). Further, $\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime}\in\mathcal{C}$ since $Z$ and $\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime}$ are topologically isomorphic by [25, Proposition 3.21 (b), p. 15] and $\mathcal{C}$ closed under topological isomorphisms. By Proposition 3.5 (a) we know that $\Delta(x)=\delta_{x}\in(\mathcal{F}(\Omega),\tau)^{\prime}$ for all $x\in\Omega$. Applying Theorem 4.5 combined with Remark 4.1 to the triple $(\Delta,\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})$, where
@@
we get that for all complete locally convex Hausdorff spaces $E$ there is a topological isomorphism $\widetilde{T}_{E}\colon\mathcal{F}(\Omega,E)_{\sigma}\to L_{b}(\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},E)$ with $\widetilde{T}_{E}(f)\circ\Delta=f$ for all $f\in\mathcal{F}(\Omega,E)_{\sigma}$, where $\widetilde{T}_{\mathbb{K}}=\mathcal{I}_{\varphi}$ by [25, Proposition 3.21 (b), p. 15]. Thus it follows from Proposition 5.8 with $\widetilde{\delta}\coloneqq\Delta$ and $Y\coloneqq\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime}$ that $(\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})\sim(\mathcal{F}(\Omega)_{\mathcal{B},\tau}^{\prime},\mathcal{I})$. We also know that $(\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})\sim(Z,\varphi)$ by [25, Proposition 3.21 (b), p. 15], which implies $(Z,\varphi)\sim(\mathcal{F}(\Omega)_{\mathcal{B},\tau}^{\prime},\mathcal{I})$. Therefore $(\mathcal{F}(\Omega),\tau)$ has a strongly unique $\mathcal{C}$ predual by Proposition 2.7. ∎

### v3

@@ 4. Linearisation of vector-valued functions
Hence, by the $\beta$-boundedness of $B$ there is some $C\geq 0$ such that $|e^{\prime}(A(x^{\prime}))|\leq C$ for all $x^{\prime}\in B\cap Y_{U}$. We deduce that the set $A(B\cap Y_{U})$ is $\sigma(E,G)$-bounded and thus bounded in $E$ because $G$ determines boundedness. We observe that $(\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime},\beta)$ is a Fréchet space by Proposition 3.5 (b) and so its linear subspace $Y_{U}$ is metrisable. We conclude that $A\colon(Y_{U},\beta_{\mid Y_{U}})\to E$ is continuous by [27, Proposition 24.10, p. 282] as the metrisable space $(Y_{U},\beta_{\mid Y_{U}})$ is bornological by [27, Proposition 24.13, p. 283]. Since $Y_{U}$ is $\beta$-dense in $\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime}$, there is a unique continuous linear extension $\widetilde{A}\colon\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime}\to E$ of $A$ by [18, 3.4.2 Theorem, p. 61–62]. Setting $F\coloneqq\chi\circ\widetilde{A}\in\mathcal{F}(\Omega,E)_{\sigma}$ by Corollary 3.6 (a) and Theorem 4.5, we observe that
@@
In particular, Theorem 4.8 is applicable to the spaces $\mathcal{F}(\Omega)=\mathcal{VC}_{P}(\Omega)$ from Example 3.7 (ii) by [25, Corollary 4.7, p. 1610–1611].
@@ 5. Equivalence and uniqueness of preduals
Moreover, (the restriction of) $T_{Z}(\widetilde{\delta})$ is a bijective map from the span of $\{\delta(x)\;|\;x\in\Omega\}$ to the span of $\{\widetilde{\delta}(x)\;|\;x\in\Omega\}$, which is dense in $Z$ by [25, Proposition 2.7, p. 1596], because $T_{Z}(\widetilde{\delta})(\delta(x))=\widetilde{\delta}(x)$ for all $x\in\Omega$. Thus $T_{Z}(\widetilde{\delta})$ has dense range.
@@
Let $z\in Z$. Since the span of $\{\widetilde{\delta}(x)\;|\;x\in\Omega\}$ is dense in $Z$ by [25, Proposition 2.7, p. 1596], there is a net $(z_{\iota})_{\iota\in I}$ converging to $z$ and where all $z_{\iota}$ can be represented as $z_{\iota}=\sum_{x\in\Omega}a_{x,\iota}\widetilde{\delta}(x)$ with finitely many non-zero $a_{x,\iota}\in\mathbb{K}$. Using that $T_{Z}(\widetilde{\delta})^{\prime\prime}\circ j_{Y}=j_{Z}\circ T_{Z}(\widetilde{\delta})$ and setting $y_{\iota}\coloneqq\sum_{x\in\Omega}a_{x,\iota}\delta(x)\in Y$ for $\iota\in I$, we get
@@
Proposition 5.1 (d) combined with Proposition 5.2 improves [1, Proposition 2.3 (c), p. 3029], [16, Theorem 3.1 (Linearization Theorem), p. 128], [29, Theorem 2.1, p. 869] and [33, Theorem 3.5, p. 19] since it shows that the Banach space $Y$ of a strong linearisation $(\delta,Y,T)$ of a Banach space $\mathcal{F}(\Omega)$ such that $T$ is an isometry is uniquely determined up to an isometric isomorphism without the need of existence of isometric isomorphisms $\widetilde{T}_{E}\colon\mathcal{F}(\Omega,E)_{\sigma}\to L(Z,E)$ such that $\widetilde{T}_{E}(f)\circ\widetilde{\delta}=f$ for all $f\in\mathcal{F}(\Omega)$ for every Banach space $E$ over $\mathbb{K}$. Further, Proposition 5.1 (d) combined with Proposition 5.2 also implies the corresponding result for the (completion of the) projective tensor product of two Banach spaces given in [35, Proposition 1.5 (Uniqueness of the Tensor Product), p. 6] and [35, Theorem 2.9, p. 22].
@@
Then it holds (a)$\Rightarrow$(b). If $\mathcal{F}(\Omega)$ is a BC-space such that $\delta_{x}\in\mathcal{F}(\Omega)^{\prime}$ for all $x\in\Omega$, then it holds (b)$\Rightarrow$(a).
@@
In particular, $(\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime},\beta)$ is a complete DF-space (see [25, Corollary 3.23, p. 1606]) and thus webbed by [18, 12.4.6 Proposition, p. 260] if $(\mathcal{F}(\Omega),\tau)$ is a Fréchet space.
@@
Using [25, Propositions 3.16, 3.17, p. 1602], Proposition 3.5 (c) and Corollary 3.6 (a), we get the following corollary for completely normable spaces $\mathcal{F}(\Omega)$.
@@
    It holds $(Z,\widetilde{T})\sim((\mathcal{F}(\Omega),\widetilde{\gamma})^{\prime},\mathcal{I})$.
@@
Now, we note that $T(\widetilde{\delta})\circ\widetilde{T}(\delta)\in L(Z)$, $\widetilde{T}(\delta)\circ T(\widetilde{\delta})\in L(Y)$ and $(T(\widetilde{\delta})\circ\widetilde{T}(\delta))(\widetilde{\delta}(x))=T(\widetilde{\delta})(\delta(x))=\widetilde{\delta}(x)$ as well as $(\widetilde{T}(\delta)\circ T(\widetilde{\delta}))(\delta(x))=\widetilde{T}(\delta)(\widetilde{\delta}(x))=\delta(x)$ for all $x\in\Omega$. Since the span of $\{\widetilde{\delta}(x)\;|\;x\in\Omega\}$ is dense in $Z$ and the span of $\{\delta(x)\;|\;x\in\Omega\}$ dense in $Y$ by [25, Proposition 2.7, p. 1596], we obtain that $T(\widetilde{\delta})$ and $\widetilde{T}(\delta)$ are topological isomorphisms with $T(\widetilde{\delta})=\widetilde{T}(\delta)^{-1}$.
@@
$\widetilde{T}_{\mathbb{K}}=(T_{Z}(\widetilde{\delta})^{t})^{-1}\circ T_{\mathbb{K}}=(T_{Z}(\widetilde{\delta})^{-1})^{t}\circ T_{\mathbb{K}}$
@@
(b)$\Rightarrow$(a) Let $(Z,\varphi)$ be a predual of $(\mathcal{F}(\Omega),\tau)$ such that $Z\in\mathcal{C}$. Due to [25, Proposition 3.21 (a), p. 1605] $(\mathcal{F}(\Omega),\tau)$ satisfies $\operatorname{(BBC)}$ and $\operatorname{(CNC)}$ for $\sigma_{\varphi}(\mathcal{F}(\Omega),Z)$ because $Z$ is complete and barrelled. By assumption for every $x\in\Omega$ there is $z_{x}\in Z$ with $\delta_{x}=\varphi(\cdot)(z_{x})$ and thus $\delta_{x}\in\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime}$ by [25, Proposition 3.21 (b), (c), p. 1605]. Hence $(\Delta,\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})$ is a strong complete barrelled linearisation of $\mathcal{F}(\Omega)$ by Corollary 3.6 (a). Further, $\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime}\in\mathcal{C}$ since $Z$ and $\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime}$ are topologically isomorphic by [25, Proposition 3.21 (b), p. 1605] and $\mathcal{C}$ closed under topological isomorphisms. By Proposition 3.5 (a) we know that $\Delta(x)=\delta_{x}\in(\mathcal{F}(\Omega),\tau)^{\prime}$ for all $x\in\Omega$. Applying Theorem 4.5 combined with Remark 4.1 to the triple $(\Delta,\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})$, where
@@
we get that for all complete locally convex Hausdorff spaces $E$ there is a topological isomorphism $\widetilde{T}_{E}\colon\mathcal{F}(\Omega,E)_{\sigma}\to L_{b}(\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},E)$ with $\widetilde{T}_{E}(f)\circ\Delta=f$ for all $f\in\mathcal{F}(\Omega,E)_{\sigma}$, where $\widetilde{T}_{\mathbb{K}}=\mathcal{I}_{\varphi}$ by [25, Proposition 3.21 (b), p. 1605]. Thus it follows from Proposition 5.8 with $\widetilde{\delta}\coloneqq\Delta$ and $Y\coloneqq\mathcal{F}(\Omega)_{\mathcal{B},\widetilde{\tau}}^{\prime}$ that $(\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})\sim(\mathcal{F}(\Omega)_{\mathcal{B},\tau}^{\prime},\mathcal{I})$. We also know that $(\mathcal{F}(\Omega)_{\mathcal{B},\sigma_{\varphi}(\mathcal{F}(\Omega),Z)}^{\prime},\mathcal{I}_{\varphi})\sim(Z,\varphi)$ by [25, Proposition 3.21 (b), p. 1605], which implies $(Z,\varphi)\sim(\mathcal{F}(\Omega)_{\mathcal{B},\tau}^{\prime},\mathcal{I})$. Therefore $(\mathcal{F}(\Omega),\tau)$ has a strongly unique $\mathcal{C}$ predual by Proposition 2.7. ∎